OpenStudy (raffle_snaffle):

The paint used to make lines on roads must reflect enough light to be clearly visible at night. Let μ denote the true average reflectometer reading for a new type of paint under consideration. A test of H0: μ = 20 versus Ha: μ > 20 will be based on a random sample of size n from a normal population distribution. What conclusion is appropriate in each of the following situations? (a) n = 14, t = 3.1, α = 0.05 (b) n = 10, t = 1.8, α = 0.01

OpenStudy (raffle_snaffle):

@jim_thompson5910

OpenStudy (raffle_snaffle):

I get the wrong answer for part a) I got P-value = 1.771 and part b) P-value = 2.821

jimthompson5910 (jim_thompson5910):

do you still have that table?

jimthompson5910 (jim_thompson5910):

in part (a), n = 14 so df = n-1 = 14-1 = 13 look in the df = 13 row only

jimthompson5910 (jim_thompson5910):

do you see any values close to t = 3.1 ?

OpenStudy (raffle_snaffle):

yes I looked at row 13 and column with 0.05 and i see 1.771

OpenStudy (raffle_snaffle):

t = 3.1 with df = 13 and alpha = 0.005

jimthompson5910 (jim_thompson5910):

the value 0.05 that you saw in the table isn't the significance level its the area to the right of the t value

OpenStudy (raffle_snaffle):

oh okay i thought it was the signifcane level. good to know

jimthompson5910 (jim_thompson5910):

well it's tied to the significance level in a way, but it's not 100% the same thing

jimthompson5910 (jim_thompson5910):

here's what the table is saying let's say we focus on the v = 1 row (ie df = 1) the first number in this row is 3.078 what does this number mean? it's under the 0.10 so the area to the right of t = 3.078 is going to be 0.10 approx. This is all for df = 1

OpenStudy (raffle_snaffle):

can you draw a pic plz?

jimthompson5910 (jim_thompson5910):